OpenStudy (anonymous):

One solution of the equation 3x2 − 16x + 5 = 0 is . What is the other solution?

OpenStudy (anonymous):

copy and paste fail

OpenStudy (anonymous):

what this is my question i need help with , iv'e been trying for the longest !

OpenStudy (anonymous):

solutions are \(\frac{1}{3}\) and 5

OpenStudy (anonymous):

\[3x^2 − 16x + 5 = 0\] \[(3x-1)(x-5)=0\] \[x=\frac{1}{3},x=5\]

OpenStudy (anonymous):

i did copy and paste am taking online class and i don't get this question ...... what s the point of the group for someone to help another person and am pretty sure people copied and paste too . THANKS

OpenStudy (anonymous):

look at your questions. it says One solution of the equation 3x2 − 16x + 5 = 0 is . there must have been something after the word "is"

OpenStudy (anonymous):

3x^2-16x+5= 0 is 1/3 . what is the other solution...

OpenStudy (mikezack123):

@satellite73 i don't think she/he ever understood you! lol